Cove’s Cakes is a local bakery. Price and cost information follows:

Price per cake $ 13.91
Variable cost per cake
Ingredients 2.26
Direct labor 1.12
Overhead (box, etc.) 0.22
Fixed cost per month $ 4,536.40


Required:
1.
Calculate Cove’s new break-even point under each of the following independent scenarios: (Round your answer to the nearest whole number.)

a. Sales price increases by $1.30 per cake.



b. Fixed costs increase by $485 per month.



c. Variable costs decrease by $0.41 per cake.



d. Sales price decreases by $0.40 per cake.



2. Assume that Cove sold 460 cakes last month. Calculate the company’s degree of operating leverage. (Do not round intermediate calculations. Round your answer to 2 decimal places.)



3. Using the degree of operating leverage calculated in Requirement 2, calculate the change in profit caused by a 10 percent increase in sales revenue. (Round your final answer to 2 decimal places (i.e. .1234 should be entered as 12.34%.))
Answer #1

Answer

  • Requirement 1

A

Price per cake

$                    15.21

B

Total variable cost per cake

$                       3.60

C = A - B

Contribution margin per cake

$                    11.61

D

Fixed Cost

$              4,536.40

E = D/C

Break Even point

391

Answer [a]

A

Price per cake

$                    13.91

B

Total variable cost per cake

$                       3.60

C = A - B

Contribution margin per cake

$                    10.31

D

Fixed Cost

$              4,721.40

E = D/C

Break Even point

458

Answer [b]

A

Price per cake

$                    13.91

B

Total variable cost per cake

$                       3.19

C = A - B

Contribution margin per cake

$                    10.72

D

Fixed Cost

$              4,536.40

E = D/C

Break Even point

423

Answer [c]

A

Price per cake

$                    13.51

B

Total variable cost per cake

$                       3.60

C = A - B

Contribution margin per cake

$                       9.91

D

Fixed Cost

$              4,536.40

E = D/C

Break Even point

458

Answer [d]

  • [2]

A

Price per cake

$                    13.91

B

Total variable cost per cake

$                       3.60

C = A - B

Contribution margin per cake

$                    10.31

D

Total cakes sold

460

E = C x D

Total Contribution margin

$              4,742.60

F

Fixed Cost

$              4,536.40

G = E - F

Net Income

$                  206.20

H = E/G

Degree of Operating Leverage

23.00

Answer

  • [3]

A

Change in Sales Revenue

10%

B

Degree of Operating Leverage

23

C = A x B

Effect on Profits

230.00%

Answer